OpenStudy (anonymous):

Solve the system using the addition method. 3x + 4y = -10 5x - 2y = 18

OpenStudy (anonymous):

what is addition method? just use an online System of equations solver.

OpenStudy (anonymous):

maybe he means add all the like terms

OpenStudy (anonymous):

it is the addition method

OpenStudy (anonymous):

it is adding some variable x or y to get rid of something. and simplify.

OpenStudy (anonymous):

im not really sure what addition method is but you can multiply -105x-2y = 18 by 2 to get -210x - 4y = 36

OpenStudy (anonymous):

then add 3x + 4y = 18 and -210x - 4y = 36 to find x